# On the Convergence of Tâtonnement for Linear Fisher Markets

### v1

@@ 4.1 Tâtonnement and price lower bounds
To prove Lemma 1, we first introduce some notations. Without loss of generality, we let $\boldsymbol{b}(j)$ be the buyer with the minimum index such that
$\boldsymbol{b}(j)\in\operatorname*{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$ (8)
@@
Now we are ready to prove Lemma 1.
Proof.
We assume there exists a $T=\min\big{\{}t\,\big{|}\,p^{t}_{j}<\tilde{p}_{j}-2m\eta\text{ for some }j\big{\}}<\infty$, otherwise we are done. In words, $T$ is the first time that some item $j$’s price goes below $\tilde{p}_{j}-2m\eta$. For each item $j$ such that $p^{T}_{j}<\tilde{p}_{j}-2m\eta$, we let $t_{j}<T$ be the last time that its price went from above $\tilde{p}_{j}$ to below $\tilde{p}_{j}$. It follows that $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq T$. Since the price decreases by at most $\eta$ per iteration, $T\geq t_{j}+2m$, and thus $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$. Note that, by definition of $T$, $p^{t_{j}}_{j^{\prime}}\geq\tilde{p}_{j^{\prime}}-2m\eta$ for all $j^{\prime}\neq j$.
By Eq. 10 and $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$, $\sum_{j^{\prime}\neq j}\sum_{t=t_{j}}^{t_{j}+2m-1}b^{t}_{\boldsymbol{b}(j)j^{\prime}}\leq 2(m-1)B_{\boldsymbol{b}(j)}$. Since from $t_{j}$ to $t_{j}+2m-1$, the total spend from buyer $\boldsymbol{b}(j)$ equals to $2mB_{\boldsymbol{b}(j)}$, $\sum_{t=t_{j}}^{t_{j}+2m-1}b^{t}_{\boldsymbol{b}(j)j}\geq 2B_{\boldsymbol{b}(j)}$. Thereby, we have
@@
Eq. 12 contradicts $p^{t_{j}+2m}_{j}\leq\tilde{p}_{j}$ and thus $T<\infty$ does not exist. By contradiction, we can conclude $p^{t}_{j}\geq\tilde{p}_{j}-2m\eta\text{ for all }j$. ∎
Next we show that the lower bound on price in Lemma 1 is tight, in terms of its order dependence on $m$.
@@
A direct consequence of Lemma 1 is that the demands are upper bounded, due to the lower bound on prices. Recall that $\sum_{i=1}^{n}B_{i}=B$.
Corollary 1 (Upper bounds of demands in the LFM).
Assume that we adjust prices in the LFM using Eq. 3, with $p^{0}\geq\tilde{p}$ and $\eta<\frac{1}{2m}\min_{j}\tilde{p}_{j}$. At any time step $t$, let $z(p^{t})\in\mathbb{R}^{m}$ be the excess demand vector and $b_{ij}^{t}$ be the spend from buyer $i$ for item $j$, then we have
Lemma 3 (Upper bounds of prices in the LFM).
Assume that we adjust prices in the LFM using Eq. 3, with $p^{0}\geq\tilde{p}$ and $\eta<\frac{1}{2m}\min_{j}\tilde{p}_{j}$, where $\tilde{p}$ is defined in Lemma 1. Then, $p^{t}_{j}\leq\left(1+\frac{\eta}{\tilde{p}_{j}-2m\eta}\right)B$ for all $t\geq 0$ and $j\in[m]$.
@@
Then, we consider the largest possible value $p^{t}_{j}$ can attain, approaching from below $B$. For any $p^{t-1}_{j}\leq B$, then $p^{t}_{j}=p^{t-1}_{j}+\eta\left(x_{j}-1\right)\leq p^{t-1}_{j}+\eta\frac{\sum_{i}b^{t}_{ij}}{p^{t-1}_{j}}\leq B+\frac{\eta B}{p^{t-1}_{j}}\leq\left(1+\frac{\eta}{\tilde{p}_{j}-2m\eta}\right)B$. Hence, the lemma follows. ∎
4.2 Quadratic growth of the dual EG objective
To guarantee a fast convergence rate for (sub)gradient descent methods, we need the objective function to satisfy some form of metric regularity condition. See Karimi et al. (2016, Appendix A) for a comprehensive comparison of different types of regularity conditions. Here, we focus on the strong convexity (SC) and quadratic growth (QG) conditions. A differentiable convex function $f$ is said to be $\mu$-strongly convex if
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\mu}{2}\lVert y-x\rVert^{2}\quad\forall\;x,y\in\text{dom }f.$ (14)
$\varphi(x)-\varphi(\Pi_{\mathcal{X}^{*}}(x))\geq\alpha\lVert x-\Pi_{\mathcal{X}^{*}}(x)\rVert^{2}\quad\forall\;x\in\text{dom }\varphi,$ (15)
where $\mathcal{X}^{*}$ is the set of the minima of $\varphi$. It is easy to see that $\mu$-strong convexity implies $\frac{\mu}{2}$-quadratic growth. Next, we show that $\varphi(p)$ in Eq. 6 satisfies QG.
Lemma 4 (Quadratic growth of $\varphi(p)$).
The convex function
satisfies the quadratic growth condition with modulus $\alpha=\min_{j}\frac{p^{*}_{j}}{2\big{(}1+\frac{\eta}{\tilde{p}_{j}-2m\eta}\big{)}^{2}B^{2}}$, where $p^{*}$ denotes the unique equilibrium price vector, and $\tilde{p}$ is defined in Lemma 1.
@@ 4.2 Quadratic growth of the dual EG objective
$h(p)=\sum_{j=1}^{m}p_{j}-\sum_{j=1}^{m}\sum_{i=1}^{n}p^{*}_{j}x^{*}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)},$
where $(p^{*},x^{*})$ is a pair of equilibrium prices and allocations. Since $h(p)$ is strictly convex, it has a unique minimum. Note that, since $\nabla_{p_{j}}h(p)=1-\frac{p^{*}_{j}}{p_{j}}\sum_{i}x^{*}_{ij}=1-\frac{p^{*}_{j}}{p_{j}}$ for all $j\in[m]$, by first-order optimality conditions we have that $p^{*}$ is the minimum of $h(p)$. Then, we have
$=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{*}_{j}x^{*}_{ij}\log{\left(\min_{k\in[m]}\frac{p_{k}}{v_{ik}}\right)}$
$\geq\sum_{j=1}^{m}p_{j}-\sum_{j=1}^{m}\sum_{i=1}^{n}p^{*}_{j}x^{*}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)}=h(p).$
Also, note that
$\varphi(p^{*})$ $=\sum_{j=1}^{m}p^{*}_{j}-\sum_{j=1}^{m}\sum_{i:x^{*}_{ij}>0}p^{*}_{j}x^{*}_{ij}\log{\left(\min_{k\in[m]}\frac{p^{*}_{k}}{v_{ik}}\right)}$
$=\sum_{j=1}^{m}p^{*}_{j}-\sum_{j=1}^{m}\sum_{i:x^{*}_{ij}>0}p^{*}_{j}x^{*}_{ij}\log{\left(\frac{p^{*}_{j}}{v_{ij}}\right)}$
$=h(p^{*}).$
Since $p$ is upper bounded by $\bar{p}$, $h$ is strongly convex with modulus $\mu=\min_{j}\frac{p^{*}_{j}}{\bar{p}_{j}^{2}}$ (see Appendix B for derivation). Since strong convexity implies the quadratic growth condition, we obtain
$h(p)-h(p^{*})\geq\min_{j}\frac{p^{*}_{j}}{2\bar{p}_{j}^{2}}\lVert p-p^{*}\rVert^{2},\quad\mbox{for all }p\leq\bar{p}.$
Since $\varphi(p)\geq h(p)$ for all $p$, and $\varphi(p^{*})=h(p^{*})$ at the unique minimum $p^{*}$,
$\varphi(p)-\varphi(p^{*})\geq h(p)-h(p^{*})\geq\min_{j}\frac{p^{*}_{j}}{2\bar{p}_{j}^{2}}\lVert p-p^{*}\rVert^{2},\quad\mbox{for all }p\leq\bar{p}.$
Therefore, $\varphi$ is a convex function satisfying the QG condition with modulus $\alpha=\min_{j}\frac{p^{*}_{j}}{2\bar{p}_{j}^{2}}$. ∎
4.3 Linear Convergence to Approximate Equilibrium
Next we show that tâtonnement converges at a linear rate, due to our results on bounded prices and quadratic growth. The proof below is typical for subgradient descent methods under these conditions, and we only include it here to show explicitly how linear convergence follows easily from our structural results on tâtonnement and the EG dual.
Theorem 1 (Linear convergence of Tâtonnement in the LFM).
$\lVert p^{t}-p^{*}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{*}\rVert^{2}+e,\quad\text{for all }t\geq 0,$ (16)
where $\alpha$ is defined in Lemma 4 and
$e=\frac{\eta G^{2}}{2\alpha}=\frac{\eta}{2\alpha}\left(\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\right)^{2}.$ (17)
Proof.
Let $g^{t}\in\partial\varphi(p^{t})$. Since $\partial\varphi(p^{t})$ is equal to the excess supply $-z^{t}$, by Corollary 1 we have $\lVert g^{t}\rVert=\lVert z^{t}\rVert\leq G:=\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m$. First we derive a descent inequality:
$\lVert p^{t+1}-p^{*}\rVert^{2}$ $=\lVert\Pi_{\mathbb{R}^{m}_{\geq 0}}\left(p^{t}-\eta g^{t}\right)-p^{*}\rVert^{2}$
$\leq\lVert p^{t}-\eta g^{t}-p^{*}\rVert^{2}$
$=\lVert p^{t}-p^{*}\rVert^{2}-2\eta\langle g^{t},p^{t}-p^{*}\rangle+\eta^{2}\lVert g^{t}\rVert^{2}$
$\leq\lVert p^{t}-p^{*}\rVert^{2}-2\eta\left(\varphi(p^{t})-\varphi(p^{*})\right)+\eta^{2}G^{2}$
$\leq\lVert p^{t}-p^{*}\rVert^{2}-2\eta\alpha\lVert p^{t}-p^{*}\rVert^{2}+\eta^{2}G^{2}$
$=(1-2\eta\alpha)\lVert p^{t}-p^{*}\rVert^{2}+\eta^{2}G^{2}.$ (18)
Equivalently, we have
$\lVert p^{t+1}-p^{*}\rVert^{2}-\frac{\eta G^{2}}{2\alpha}\leq(1-2\eta\alpha)\left(\lVert p^{t}-p^{*}\rVert^{2}-\frac{\eta G^{2}}{2\alpha}\right).$ (19)
By induction, we obtain that $\lVert p^{t}-p^{*}\rVert^{2}-e\leq(1-2\eta\alpha)^{t}\left(\lVert p^{0}-p^{*}\rVert^{2}-e\right)\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{*}\rVert^{2}$ where $e=\frac{\eta G^{2}}{2\alpha}$. ∎
Eq. 17 shows that the price vector $p^{t}$ converges linearly to an $\epsilon$ neighborhood around the equilibrium price vector $p^{*}$, where the size of the neighborhood is determined by the stepsize $\eta$. Suppose instead that we want to get arbitrarily close to $p^{*}$. Next, we show a corrolary of Eq. 17 for this case.
Corollary 2 (Time complexity of Tâtonnement in the LFM).
For a given $\epsilon>0$, let $\eta\leq\min\left\{\frac{\min_{j}\tilde{p}_{j}}{4m},\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}\right\}$. Then, starting from any $p^{0}\geq\tilde{p}$ such that $\sum_{j=1}^{m}p^{0}_{j}=B$, tâtonnement with stepsize $\eta$ generates a price vector $p$ such that $\lVert p-p^{*}\rVert^{2}\leq\epsilon$ in $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$ iterations. If $\epsilon\geq\frac{9B^{2}\min_{j}\tilde{p}_{j}}{8m\min_{j}p^{*}_{j}}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}$, the time complexity is on the order of $\mathcal{O}(\log{\frac{1}{\epsilon}})$.
Proof.
@@ 4.3 Linear Convergence to Approximate Equilibrium
Since $\eta\leq\frac{\min_{j^{\prime}}\tilde{p}_{j^{\prime}}}{4m}\leq\frac{\tilde{p}_{j}}{4m}$ for all $j\in[m]$, we have $\frac{\eta}{\tilde{p}_{j}-2m\eta}\leq\frac{1}{2m}$ and thus $\alpha\geq\frac{2\min_{j}p^{*}_{j}}{9B^{2}}$. It follows that
$e=\frac{\eta}{2\alpha}\left(\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\right)^{2}\leq\frac{9\eta B^{2}}{4\min_{j}p^{*}_{j}}\left(\frac{2B}{\min_{j}\tilde{p}_{j}}+m\right)^{2}\leq\frac{\epsilon}{2},$ (20)
@@
where the inequality follows by $\eta\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$ and the last equality follows by $\eta\leq\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}$. After $t=\frac{1}{2\eta\alpha}\log{\frac{8B^{2}}{\epsilon}}$ iterations, we have
$t\log{(1-2\eta\alpha)}+\log{\lVert p^{0}-p^{*}\rVert^{2}}\leq-2\eta\alpha\cdot t+\log{4B^{2}}=-\log{\frac{8B^{2}}{\epsilon}}+\log{4B^{2}}=\log{\frac{\epsilon}{2}},$ (21)
@@
where the inequality follows because $\log(1+x)\leq x,\;\forall\,x>-1$ and $\lVert p^{0}-p^{*}\rVert\leq 2B$. Thus, by Eq. 17 $\lVert p^{t}-p^{*}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{*}\rVert^{2}+e\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$.
@@
Since $\alpha\geq\frac{2\min_{j}p^{*}_{j}}{9B^{2}}$, $t\leq\frac{9B^{2}}{4\eta\min_{j}p^{*}_{j}}\log(\frac{8B^{2}}{\epsilon})$. If $\epsilon$ is small such that $\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$, $\eta$ is of the order $\mathcal{O}(\epsilon)$, hence the time complexity is of the order $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$. Otherwise, $\eta$ is independent of $\epsilon$, hence the time complexity is of the order $\mathcal{O}(\log{\frac{1}{\epsilon}})$. ∎

### v2

@@ Appendix B Proofs in Section 4
    To prove Lemma 1 $(i)$, we first introduce some notations. Without loss of generality, we let $\boldsymbol{b}(j)$ be the buyer with the minimum index such that
    $\boldsymbol{b}(j)\in\operatorname\textup{*}{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$ (15)
@@
    Now we are ready to prove Lemma 1. We assume there exists a $T=\min\big{\{}t\,\big{|}\,p^{t}_{j}<\tilde{p}_{j}-2m\eta\text{ for some }j\big{\}}<\infty$, otherwise we are done. In words, $T$ is the first time that some item $j$’s price goes below $\tilde{p}_{j}-2m\eta$. For each item $j$ such that $p^{T}_{j}<\tilde{p}_{j}-2m\eta$, we let $t_{j}<T$ be the last time that its price went from above $\tilde{p}_{j}$ to below $\tilde{p}_{j}$. It follows that $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq T$. Since the price decreases by at most $\eta$ per iteration, $T\geq t_{j}+2m$, and thus $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$. Note that, by definition of $T$, $p^{t_{j}}_{j^{\prime}}\geq\tilde{p}_{j^{\prime}}-2m\eta$ for all $j^{\prime}\neq j$.
    By Eq. 17 and $p^{t}_{j}\leq\tilde{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$, $\sum_{j^{\prime}\neq j}\sum_{t=t_{j}}^{t_{j}+2m-1}b^{t}_{\boldsymbol{b}(j)j^{\prime}}\leq 2(m-1)B_{\boldsymbol{b}(j)}$. Since from $t_{j}$ to $t_{j}+2m-1$, the total spend from buyer $\boldsymbol{b}(j)$ equals $2mB_{\boldsymbol{b}(j)}$, $\sum_{t=t_{j}}^{t_{j}+2m-1}b^{t}_{\boldsymbol{b}(j)j}\geq 2B_{\boldsymbol{b}(j)}$. Thereby, we have
@@
    Eq. 19 contradicts $p^{t_{j}+2m}_{j}\leq\tilde{p}_{j}$ and thus $T<\infty$ does not exist. By contradiction, we can conclude $p^{t}_{j}\geq\tilde{p}_{j}-2m\eta\text{ for all }j$. This completes the proof of Lemma 1 $(i)$.
  • (ii)
    A direct consequence of Lemma 1 $(i)$ is that the demands are upper bounded, due to the lower bound on prices. Recall that $\sum_{i=1}^{n}B_{i}=B$. At any time step $t$, let $z(p^{t})\in\mathbb{R}^{m}$ be the excess demand vector, then we have
  • (iii)
    Then, we consider the largest possible value $p^{t}_{j}$ can attain, approaching from below $B$. For any $p^{t-1}_{j}\leq B$, then $p^{t}_{j}=p^{t-1}_{j}+\eta\left(x_{j}-1\right)\leq p^{t-1}_{j}+\eta\frac{\sum_{i}b^{t}_{ij}}{p^{t-1}_{j}}\leq B+\frac{\eta B}{p^{t-1}_{j}}\leq\left(1+\frac{\eta}{\tilde{p}_{j}-2m\eta}\right)B$. Hence, Lemma 1 $(iii)$ follows.
@@
Strong convexity of $h$.
Here, we compute the Hessian matrix of $h(p)$ and then determine the strong convexity modulus of $h(p)$. First, note that
$h(p)=\sum_{j}p_{j}-\sum_{j}\sum_{i}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)}$ $=\sum_{j}p_{j}-\sum_{j}p^{\textup{*}}_{j}\Big{(}\sum_{i}x^{\textup{*}}_{ij}\log{p_{j}}-\sum_{i}x^{\textup{*}}_{ij}\log{v_{ij}}\Big{)}$
$=\sum_{j}p_{j}-\sum_{j}p^{\textup{*}}_{j}\Big{(}\log{p_{j}}-\sum_{i}x^{\textup{*}}_{ij}\log{v_{ij}}\Big{)}$
$=\sum_{j}(p_{j}-p^{\textup{*}}_{j}\log{p_{j}})-\sum_{i}\sum_{j}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}\log{v_{ij}}.$
It follows that the Hessian matrix
$\nabla^{2}h(p)=\begin{bmatrix}\frac{p^{\textup{*}}_{1}}{p_{1}^{2}}&0&\cdots&0\\ 0&\frac{p^{\textup{*}}_{2}}{p_{2}^{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\frac{p^{\textup{*}}_{m}}{p_{m}^{2}}\end{bmatrix}\succeq\min_{j}\frac{p^{\textup{*}}_{j}}{\bar{p}_{j}^{2}}\;\textbf{I},$ (21)
where I denotes the $m$-dimensional identity matrix. Equivalently, $h(p)$ is strongly convex with modulus $\mu=\min_{j}\frac{p^{\textup{*}}_{j}}{\bar{p}_{j}^{2}}$. The strong convexity modulus of $h^{q}(p)$ can be computed in the same way.
Proofs of Theorem 1 and Corollary 1.
Theorem 1.
$\lVert p^{t}-p^{\textup{*}}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{\textup{*}}\rVert^{2}+e,\quad\text{for all }t\geq 0,$
where $\alpha$ is defined in Lemma 3 and
$e=\frac{\eta G^{2}}{2\alpha}=\frac{\eta}{2\alpha}\Big{(}\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\Big{)}^{2}.$
@@
Let $g^{t}\in\partial\varphi(p^{t})$. Since $\partial\varphi(p^{t})$ is equal to the excess supply $-z^{t}$, by Lemma 1 $(ii)$, we have $\lVert g^{t}\rVert=\lVert z^{t}\rVert\leq G:=\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m$. First we derive a descent inequality:
$\lVert p^{t+1}-p^{\textup{*}}\rVert^{2}$ $=\lVert\Pi_{\mathbb{R}^{m}_{\geq 0}}\left(p^{t}-\eta g^{t}\right)-p^{\textup{*}}\rVert^{2}$
$\leq\lVert p^{t}-\eta g^{t}-p^{\textup{*}}\rVert^{2}$
$=\lVert p^{t}-p^{\textup{*}}\rVert^{2}-2\eta\langle g^{t},p^{t}-p^{\textup{*}}\rangle+\eta^{2}\lVert g^{t}\rVert^{2}$
$\leq\lVert p^{t}-p^{\textup{*}}\rVert^{2}-2\eta\left(\varphi(p^{t})-\varphi(p^{\textup{*}})\right)+\eta^{2}G^{2}$
$\leq\lVert p^{t}-p^{\textup{*}}\rVert^{2}-2\eta\alpha\lVert p^{t}-p^{\textup{*}}\rVert^{2}+\eta^{2}G^{2}$
$=(1-2\eta\alpha)\lVert p^{t}-p^{\textup{*}}\rVert^{2}+\eta^{2}G^{2}.$ (22)
Equivalently, we have
$\lVert p^{t+1}-p^{\textup{*}}\rVert^{2}-\frac{\eta G^{2}}{2\alpha}\leq(1-2\eta\alpha)\left(\lVert p^{t}-p^{\textup{*}}\rVert^{2}-\frac{\eta G^{2}}{2\alpha}\right).$ (23)
By induction, we obtain that $\lVert p^{t}-p^{\textup{*}}\rVert^{2}-e\leq(1-2\eta\alpha)^{t}\left(\lVert p^{0}-p^{\textup{*}}\rVert^{2}-e\right)\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{\textup{*}}\rVert^{2}$ where $e=\frac{\eta G^{2}}{2\alpha}$. ∎
Corollary 1.
For a given $\epsilon>0$, let $\eta\leq\min\left\{\frac{\min_{j}\tilde{p}_{j}}{4m},\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}\right\}$. Then, starting from any $p^{0}\geq\tilde{p}$ such that $\sum_{j=1}^{m}p^{0}_{j}=B$, tâtonnement with stepsize $\eta$ generates a price vector $p$ such that $\lVert p-p^{\textup{*}}\rVert^{2}\leq\epsilon$ in $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$ iterations. If $\epsilon\geq\frac{9B^{2}\min_{j}\tilde{p}_{j}}{8m\min_{j}p^{\textup{*}}_{j}}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}$, the time complexity is on the order of $\mathcal{O}(\log{\frac{1}{\epsilon}})$.
@@
Since $\eta\leq\frac{\min_{j^{\prime}}\tilde{p}_{j^{\prime}}}{4m}\leq\frac{\tilde{p}_{j}}{4m}$ for all $j\in[m]$, we have $\frac{\eta}{\tilde{p}_{j}-2m\eta}\leq\frac{1}{2m}$ and thus $\alpha\geq\frac{2\min_{j}p^{\textup{*}}_{j}}{9B^{2}}$. It follows that
$e=\frac{\eta}{2\alpha}\left(\frac{B}{\min_{j}\tilde{p}_{j}-2m\eta}+m\right)^{2}\leq\frac{9\eta B^{2}}{4\min_{j}p^{\textup{*}}_{j}}\left(\frac{2B}{\min_{j}\tilde{p}_{j}}+m\right)^{2}\leq\frac{\epsilon}{2},$ (24)
@@
where the inequality follows by $\eta\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$ and the last equality follows by $\eta\leq\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}$. After $t=\frac{1}{2\eta\alpha}\log{\frac{8B^{2}}{\epsilon}}$ iterations, we have
$t\log{(1-2\eta\alpha)}+\log{\lVert p^{0}-p^{\textup{*}}\rVert^{2}}\leq-2\eta\alpha\cdot t+\log{4B^{2}}=-\log{\frac{8B^{2}}{\epsilon}}+\log{4B^{2}}=\log{\frac{\epsilon}{2}},$ (25)
@@
where the inequality follows because $\log(1+x)\leq x,\;\forall\,x>-1$ and $\lVert p^{0}-p^{\textup{*}}\rVert\leq 2B$. Thus, by Theorem 1 $\lVert p^{t}-p^{\textup{*}}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{\textup{*}}\rVert^{2}+e\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$.
@@
Since $\alpha\geq\frac{2\min_{j}p^{\textup{*}}_{j}}{9B^{2}}$, $t\leq\frac{9B^{2}}{4\eta\min_{j}p^{\textup{*}}_{j}}\log(\frac{8B^{2}}{\epsilon})$. If $\epsilon$ is small such that $\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\tilde{p}_{j}}+m)^{2}}\leq\frac{\min_{j}\tilde{p}_{j}}{4m}$, $\eta$ is of the order $\mathcal{O}(\epsilon)$, hence the time complexity is of the order $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$. Otherwise, $\eta$ is independent of $\epsilon$, hence the time complexity is of the order $\mathcal{O}(\log{\frac{1}{\epsilon}})$. ∎
Appendix C More Detailed Discussion on the LFM
Approximate equilibrium implied by Theorem 1.